# Robust, fast, and efficient formation of stable tetratomic molecules from ultracold atoms via generalized stimulated Raman exact passage

### v3

@@ II.2 Result and Discussion
Figure 2: (Color online) Rabi frequencies (left column) and populations (left column) as a function of $t$. Adopted Parameters: $\eta=0.96,\epsilon=0.2$.
To demonstrate the validity of our procedure, we are going to carry out numerical calculations on population dynamics by employing the standard fourth-order Runge-Kutta method.
@@
To better illustrate the advantages of our protocol, we simulated its performance and compared it with that of the conventional STIRAP protocol. This study is feasible, because the transfer process we consider can be viewed as an abstract three-level model. In our simulations, the STIRAP process was implemented using a pair of Gaussian pulses Meng et al. (2008),
@@
where $T$, $t_{\mathrm{P,S}}$ and $\Omega^{0}_{\mathrm{P,S}}$ denote the width, peak times and peak strengths of the Gaussian pulses, respectively. The Stokes pulse $\Omega_{\mathrm{S}}$ is applied first to induce the transition $|m\rangle\leftrightarrow|t\rangle$ and is followed by the pump pulse $\Omega_{\mathrm{P}}$, which induces the transition $|a\rangle\leftrightarrow|m\rangle$.
Figure 3: (Color online) (a) Rabi frequencies as functions of $t$ under the generalized STIREP and the STIRAP. $\chi=8.4437$; (b) Conversion efficiency as a function of $\eta_{1}$ and $\eta_{2}$ under the (upper surface) generalized STIREP and (lower surface) STIRAP protocols; (c) Conversion efficiency as a function of $T$ under the generalized STIREP; Parameters: $\gamma_{m}=10^{-1}/T$, $\gamma_{t}=10^{-4}/T$.
@@
To ensure a fair comparison, both schemes are carried out with the same evolution time, peak times, and peak strengths. According to Fig. 3(a), we set $t_{\mathrm{P}}\!=\!1.727T$, $t_{\mathrm{S}}\!=\!0.006T$, $\Omega^{0}_{\mathrm{P}}=27.7204/T$, and $\Omega^{0}_{\mathrm{S}}=3.283/T$, respectively. Meanwhile, the robustness of both schemes against Rabi frequency errors will be assessed. To this end, we systematically modify the amplitudes of the original Rabi frequencies while preserving their waveform shapes, with specific adjustments defined as $\Omega_{\mathrm{1,P}}\rightarrow(1-\eta_{1})\Omega_{\mathrm{1,P}}$ and $\Omega_{\mathrm{2,S}}\rightarrow(1-\eta_{2})\Omega_{\mathrm{2,S}}$, respectively.
@@
Figure. 3(a) displays the time sequence of the Rabi frequencies corresponding to our protocol and STIRAP. It is obvious that, like the Gaussian waveform, the fields used in our protocol are still sufficiently smooth. We plot in Fig. 3(b) the conversion efficiency as a function of $\eta_{1}$ and $\eta_{2}$. Here, the upper surface corresponds to our protocol, whereas the lower surface corresponds to STIRAP. As it is observed, our protocol always performs better than STIRAP throughout the whole range of error parameters. In addition, we stress that our protocol allows for faster control over the conversion process. We demonstrate this fact by plotting a curve of the conversion efficiency of our protocol as a function of the evolution time $T$, as depicted in Fig. 3(c). As it turns out, the higher conversion efficiency can be achieved by shortening the evolution time. This intriguing phenomenon aligns with the core spirit of STA.
@@
Further studies can reveal more properties of our protocol. In Fig. 4(a), by fixing $\Omega_{2}$ at its ideal value, we present a contour plot of transfer efficiency against the evolution time deviation $\delta T$ and the Rabi frequency deviation $\delta\Omega_{1}$. As a comparison, we study in Fig. 4(b) transfer efficiency against $\delta T$ and $\delta\Omega_{2}$, with $\Omega_{1}$ fixed at its ideal value. It should be noted that the values marked by the white dashed lines in these figures are equal to the ideal value under the working condition of no deviations; this value correspond exactly to the end value of the red curve in Fig. 2(b), and its specific value is $0.9899$. It is observed that even for deviations of $|\delta T/T|\!=\!|\delta\Omega_{1}/\Omega_{1}|\!=\!|\delta\Omega_{2}/\Omega_{2}|\!=\!10\%$, the conversion efficiencies remain remarkably close to the ideal value. The results further demonstrate that the C-STIREP exhibits high robustness with respect to imperfections in operation.
@@
The above calculations demonstrate that the robustness and fast conversion of our protocol facilitate the preparation of ultracold tetratomic molecules. Nevertheless, it should be stressed that the resulting molecules are usually energetically unstable Kuznetsova et al. (2009), thereby making it highly necessary from a practical standpoint to rapidly transfer them to deeply-bound vibrational states. To this end, we suggest a five-state C-STIREP technique, whose details will be further analyzed and presented in the following section.
@@ III.1 Model and Method
In this section, our goal is to generalize the STIREP into multi-state system, with the motivation to transfer the weakly-bound molecules into a sufficiently stable state. Inspired by the double-STIRAP transfer scheme Danzl et al. (2010); Kuznetsova et al. (2008), we restrict ourselves to a five-level chainwise-coupled molecular condensate system. As presented in Fig. 5, all five molecular states are coupled in a distorted M-type configuration,
@@
in which four fields with Rabi frequencies $\Omega_{i}(i=1,2,3,4)$ are applied to move the molecules from a high vibrational level of the ground electronic state $|c\rangle$ (now relabeled as $|1\rangle$) to a sufficiently stable vibrational level of the ground electronic state $|5\rangle$. In the entire transfer process, three intermediate states are introduced to act as the “bridge”, where $|3\rangle$ is regarded as an intermediate vibrational level of ground electronic state, $|2\rangle$ and $|4\rangle$ represent two vibrational levels of the excited electronic state. To render this problem tractable, we shall analyze this problem under the collisionless limit, as done in prior Refs. Pu et al. (2007); Jing et al. (2008); Qian et al. (2010); Graefe et al. (2015). Accordingly, the Hamiltonian of the system can be described in the following form:
$\begin{aligned} \hat{\mathcal{H}}\!=\!\sum_{i=2}^{5}\Delta_{i}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{i}\!+\!\sum_{i=1}^{4}(\Omega_{i}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{i+1}\!+\!\mathrm{H.c.}).\end{aligned}$ (9)
@@
$\begin{aligned} i\dot{\psi}_{1}&\!=\!\Omega_{1}\psi_{2},\\ i\dot{\psi}_{2}&\!=\!\Omega_{1}\psi_{1}\!+\!\Delta_{1}\psi_{2}\!+\!\Omega_{2}\psi_{3},\\ i\dot{\psi}_{3}&\!=\!\Omega_{2}\psi_{2}\!+\!\Omega_{3}\psi_{4},\\ i\dot{\psi}_{4}&\!=\!\Omega_{3}\psi_{3}\!+\!\Delta_{2}\psi_{4}\!+\!\Omega_{4}\psi_{5},\\ i\dot{\psi}_{5}&\!=\!\Omega_{4}\psi_{4}.\end{aligned}$ (10)
@@
First, we postulate that $\Delta$ is the largest evolution frequency Lugiato et al. (1984); Mackie and Phou (2010); Liu et al. (2019b), $\Delta\!\gg\!\Omega_{i}(i\!=\!1,2,3,4)$. In this limit, adiabatic elimination (AE) enables the derivation of an effective three-state model, which takes the form of
@@
In which, the effective Rabi frequencies are $\omega_{13}=\!-\!\frac{\Omega_{1}\Omega_{2}}{\Delta}$ and $\omega_{35}=\!-\!\frac{\Omega_{3}\Omega_{4}}{\Delta}$; the effective detunings are $\varepsilon_{11}=\!-\!\frac{\Omega_{1}^{2}}{\Delta},\varepsilon_{22}=\!-\!\frac{\Omega_{2}^{2}+\Omega_{3}^{2}}{\Delta}$ and $\varepsilon_{33}=\!-\!\frac{\Omega_{4}^{2}}{\Delta}$, respectively. To continue, let us set $\varepsilon_{11}=\varepsilon_{22}=\varepsilon_{33}=\varepsilon_{0}$, from which the condition for the generalized resonance locking follows,
@@
By further performing the simple transformation $\psi_{j}\!=\!\psi^{{}^{\prime}}_{j}e^{-i\varepsilon_{0}t}(j\!=\!1,3,5\!)$, we can obtain
@@
$\begin{aligned} \omega_{13}\!=-\!\frac{\Omega_{2}\sqrt{\Omega_{2}^{2}\!+\!\Omega_{3}^{2}}}{\Delta},\quad\omega_{35}\!=-\!\frac{\Omega_{3}\sqrt{\Omega_{3}^{2}\!+\!\Omega_{2}^{2}}}{\Delta}.\end{aligned}$ (14)
@@
$\begin{aligned} &\psi^{{}^{\prime}}_{1}\!=\!\cos\phi\cos\theta,\\ &\psi^{{}^{\prime}}_{3}\!=\!-i\sin\phi,\\ &\psi^{{}^{\prime}}_{5}\!=\!-\cos\phi\sin\theta,\end{aligned}$ (15)
@@
$\begin{aligned} \omega_{13}&=\dot{\theta}\cot\phi\sin\theta\!+\!\dot{\phi}\cos\theta,\\ \omega_{35}&=\dot{\theta}\cot\phi\cos\theta\!-\!\dot{\phi}\sin\theta.\end{aligned}$ (16)
@@
Now we are ready to design the modified Rabi frequencies for the original system. Like Eq. (14), we impose the following form on the two middle Rabi frequencies:
$\begin{aligned} \omega_{13}=-\frac{\tilde{\Omega}_{2}\sqrt{\tilde{\Omega}_{2}^{2}\!+\!\tilde{\Omega}_{3}^{2}}}{\Delta},\quad\omega_{35}=-\frac{\tilde{\Omega}_{3}\sqrt{\tilde{\Omega}_{2}^{2}\!+\!\tilde{\Omega}_{3}^{2}}}{\Delta}.\end{aligned}$ (17)
@@
$\begin{aligned} \tilde{\Omega}_{1}=\tilde{\Omega}_{4}=\sqrt{\tilde{\Omega}_{2}^{2}+\tilde{\Omega}_{3}^{2}}.\end{aligned}$ (18)
@@
$\begin{aligned} \tilde{\Omega}_{2}&\!=\!\left(\frac{\Delta^{2}\omega^{4}_{13}}{\omega^{2}_{13}\!+\!\omega^{2}_{35}}\right)^{\frac{1}{4}},\quad\tilde{\Omega}_{3}\!=\!\left(\frac{\Delta^{2}\omega^{4}_{35}}{\omega^{2}_{13}\!+\!\omega^{2}_{35}}\right)^{\frac{1}{4}},\\ \tilde{\Omega}_{1}&\!=\!\tilde{\Omega}_{4}\!=\!\left[\Delta^{2}(\omega^{2}_{13}\!+\!\omega^{2}_{35})\right]^{\frac{1}{4}}.\end{aligned}$ (19)
With the newly designed Rabi frequencies, we thereby develop a five-state C-STIREP technique. It should be noted that, to ensure the validity of this method, the AE condition $\Delta\!\gg\!\tilde{\Omega}_{i}(i\!=\!1,2,3,4)$ should still be satisfied Li and Chen (2016).
@@ III.2 Result and Discussion
Now we numerically investigate the coherent population dynamics of the C-STIREP method. Unlike the above (7), we demonstrate that the enhanced performance of C-STIREP can also be achieved by using other pulse shapes. As an example, we consider
@@
Once $\phi$ and $\theta$ are determined, the effective Rabi frequencies can be found according to Eq. (16).
@@
Figure. 6(a) illustrates the Rabi frequencies required to achieve the desired population transfer result. From this figure, it can be observed that the first and fourth Rabi frequencies not only exhibit the same temporal profile and equal magnitudes, but also straddle the middle two ones. They feature near-Gaussian profiles with sufficient smoothness, thus posing no actual implementation challenges Du et al. (2016); Zheng et al. (2022); Norambuena et al. (2024). The time evolution results from the initial state $|1\rangle$ to the target state $|5\rangle$ is displayed in Fig. 6(b). It can be observed that a complete state transfer can be achieved by employing four modified fields. A high transfer efficiency is helpful to minimize the loss in phase-space density when an ensemble of weakly-bound molecules is transferred to the ground state. In addition, we note that, throughout the entire transfer process, the intermediate state $|3\rangle$ receives some amount of transient population, while the populations of the two excited states are negligible. This process does not rely on the existence of dark states; instead, we utilize the AE protocol to achieve the stable decoupling of two excited states from dynamics. Thus, although these two states are prone to decaying out of the system, the transfer process remains scarcely swayed by them.
@@
Figure 7: (Color online) (a) The transient population of state $|3\rangle$ as a function of $\beta$, where the other parameters are consistent with those in Fig. 6; (b) The populations of state $|3\rangle$ and state $|5\rangle$ as a function of $T$, where $\beta$ is fixed at $0.1\pi$, and all other parameters are consistent with those in Fig. 6.
@@
Once a desired population transfer from $|1\rangle$ to $|5\rangle$ is achieved, we would like to impose further conditions of minimizing the amount of transient population that is in $|3\rangle$ throughout the process. Since the dynamics of the five-state system is equivalent to that of the reduced three-state one, the transient population of the intermediate state $|3\rangle$ can be accurately derived as $\sin^{2}\phi$. On this basis, we may attempt to reduce the value of $\phi$ to diminish the population of this state. We plot in Fig. 7(a) the transient population of state $|3\rangle$ as a function of $\beta$; please note that all the other parameters are maintained the same as those in Fig. 6. Clearly, the population of the intermediate state $|3\rangle$ witnesses a substantial decrease as the value of $\beta$ is gradually reduced. However, the population of this state never equals 0 since the limit of $\beta\rightarrow 0$ cannot be achieved in practical scenarios [see Eqs. (19) and (20)]. From the preceding analysis, it can be concluded that the five-state transfer scheme offers an additional benefit over the three-state counterpart: the former only requires us to control the transient population of the intermediate state $|3\rangle$, while the latter necessitates the control of the short-lived excited state. Given that the intermediate state $|3\rangle$ typically exhibits a sufficiently long lifetime, even a small amount of transient population within this state exerts a negligible influence on the final transfer efficiency; in contrast, the transfer efficiency of the three-state scheme is largely constrained by the characteristics of its intermediate short-lived excited state Ciamei et al. (2017); Li et al. (2017).
@@
Figure 8: (Color online) (a) Contour plot showing transfer efficiency versus $\delta\tilde{\Omega}_{1}$ and $\delta\tilde{\Omega}_{4}$, where $\Omega_{2}$ and $\Omega_{3}$ are set to their ideal values without any deviation; (b) Contour plot showing transfer efficiency versus $\delta\tilde{\Omega}_{2}$ and $\delta\tilde{\Omega}_{3}$, where $\Omega_{2}$ and $\Omega_{3}$ passively vary according to condition (18). The parameter deviations are implemented on the basis of the adopted parameters in Fig. 7(b).
@@
Furthermore, we point out that the transient population of the intermediate state $|3\rangle$ can be further reduced by leveraging the speed advantage of C-STIREP. To verify this finding, we plot in Fig. 7(b) the variation of populations of state $|3\rangle$ and state $|5\rangle$ with the evolution time $T$. As can be seen from this figure, the proposed method is capable of significantly shortening the population transfer time without deteriorating the transfer efficiency. This interesting feature mitigates the interaction duration between the system and its surrounding environment, thereby enhancing the transfer efficiency for practical applications.
Finally, we are going to perform the targeted numerical simulations to verify the necessity of Condition (18). Figure. 8(a) shows a contour plot of transfer efficiency against the Rabi frequency deviations $\delta\tilde{\Omega}_{1}$ and $\delta\tilde{\Omega}_{4}$ when $\tilde{\Omega}_{2}$ and $\tilde{\Omega}_{3}$ are fixed at their ideal values. Clearly, $\delta\!=\!0$ implies that an ideal operating condition is achieved; $\delta\!\neq\!0$ indicates that the Condition (18) is violated. According to the results shown in this figure, we can find that any deviation from the ideal condition will result in a significant drop in efficiency. In particular, when the deviation is relatively large, the transfer efficiency will suffer a loss of around $10\%$. Except for this, we can observe that the deviations $\tilde{\Omega}_{1}$ and $\tilde{\Omega}_{4}$ exert symmetric effects on the transfer efficiency. This confirms that this condition is an indispensable prerequisite for obtaining high population transfer efficiency. The transfer efficiency can remain at a high level if and only if $\tilde{\Omega}_{1}\!=\!\tilde{\Omega}_{4}$ and they should exactly match the ideal value.

### v4

@@ II.2 Result and Discussion
Figure 2: (Color online) Rabi frequencies (top panel) and populations (bottom panel) as a function of $t$. Adopted Parameters: $\eta=0.96,\epsilon=0.2,\gamma_{\mathrm{m,t}}=0$.
To demonstrate the validity of our procedure, we are going to carry out numerical calculations for population dynamics using Eqs. (6), (8) and (9), with the standard fourth-order Runge-Kutta method.
@@
In addition, we stress that our protocol allows for faster control over the conversion process at the same conversion efficiency. We demonstrate this fact by plotting a curve of the conversion efficiency of our protocol as a function of the evolution time $T$, as depicted by the solid red line in Fig. 3(a). This intriguing phenomenon aligns with the core spirit of STA. Owing to the exact dynamics, we find that although the maximum transient population of the intermediate triatomic molecular state is independent of the value of $T$, the total transient population of this state over the entire process can be reduced by shortening the evolution time $T$. At the same time, given that $\epsilon$ is a key parameter governing the transient population of the intermediate triatomic state, we investigate the maximum transient population of the intermediate triatomic molecules and the final population of tetratomic molecules as a function of parameter $\epsilon$ in Fig. 3(b). The results presented in this figure show that decreasing the value of parameter $\epsilon$ not only helps to improve the conversion efficiency $\small{4}|\psi_{\mathrm{t}}\small{(t\!=\!t_{f})}|^{2}$, but also effectively suppresses the transient population of intermediate triatomic molecules. From the analysis of this figure, we can thus conclude that the decay of the intermediate triatomic state can be minimized via two ways, namely the reasonable selection of a small value of the parameter $\epsilon$, and the shortening of the evolution time.
Figure 3: (Color online) The maximum transient population of intermediate triatomic molecules and the final population of tetratomic molecules as a function of parameter $\epsilon$. All other parameters are consistent with those in Fig. 2.
To better illustrate the advantages of our protocol, we simulated its performance and compared it with that of the conventional STIRAP protocol. This study is feasible, because the transfer process we consider can be viewed as an abstract three-level model. In our simulations, the STIRAP process was implemented using a pair of Gaussian pulses,
@@
where $T$, $t_{\mathrm{P,S}}$ and $\Omega^{0}_{\mathrm{P,S}}$ denote the width, peak times and peak strengths of the Gaussian pulses, respectively. The Stokes pulse $\Omega_{\mathrm{S}}$ is applied first to induce the transition $|\mathrm{m}\rangle\leftrightarrow|\mathrm{t}\rangle$ and is followed by the pump pulse $\Omega_{\mathrm{P}}$, which induces the transition $|\mathrm{a}\rangle\leftrightarrow|\mathrm{m}\rangle$.
Figure 4: (Color online) (a) Rabi frequencies as functions of $t$ under the generalized STIREP and the STIRAP. Adopted Parameters: $\eta=0.96,\epsilon=0.2$; (b) Conversion efficiency as a function of $\eta_{1}$ and $\eta_{2}$ under the (upper surface) generalized STIREP and (lower surface) STIRAP protocols.
@@
To ensure a fair comparison, both schemes are carried out with the same evolution time, peak times, and peak strengths. According to Fig. 2(a), where the curves are determined by Eq. (8) and Eq. (9) [also presented in Fig. 4(a)], we accordingly set the amplitudes of the Gaussian envelopes to $\Omega^{0}_{\mathrm{P}}=27.7204/T$and $\Omega^{0}_{\mathrm{S}}=3.283/T$, and the peak times to $t_{\mathrm{P}}\!=\!1.727T$and $t_{\mathrm{S}}\!=\!0.006T$. Specifically, the amplitudes of the Gaussian envelopes are directly obtained from the vertical coordinates corresponding to the peaks of the curves, while the peak times are directly derived from the horizontal coordinates corresponding to these peaks. Meanwhile, the robustness of both schemes against Rabi frequency errors will be assessed. To this end, we systematically modify the amplitudes of the original Rabi frequencies while preserving their waveform shapes, with specific adjustments defined as $\Omega_{\mathrm{1,P}}\rightarrow(1\!-\!\eta_{1})\Omega_{\mathrm{1,P}}$ and $\Omega_{\mathrm{2,S}}\rightarrow(1\!-\!\eta_{2})\Omega_{\mathrm{2,S}}$, respectively.
@@
Figure. 4(a) displays the time dependence of the Rabi frequencies corresponding to our protocol and STIRAP. It is obvious that the fields used in our protocol exhibit profiles with sufficient smoothness. We plot in Fig. 4(b) the conversion efficiency as a function of $\eta_{1}$ and $\eta_{2}$. Here, the upper surface corresponds to our protocol, whereas the lower surface corresponds to STIRAP. As it is observed, our protocol always performs better than STIRAP.
@@
Further studies can reveal more properties of our protocol. In Fig. 5(a), by fixing $\Omega_{2}$ at its ideal value, we present a contour plot of transfer efficiency against the evolution time deviation $\delta T$ and the Rabi frequency deviation $\delta\Omega_{1}$. For any physical quantity $X$, its deviation is explicitly defined in Appendix B. As a comparison, we study in Fig. 5(b) transfer efficiency against $\delta T$ and $\delta\Omega_{2}$, with $\Omega_{1}$ fixed at its ideal value. It should be noted that the values marked by the white dashed lines in these figures are equal to the ideal value under the working condition of no deviations; this value correspond exactly to the end value of the red curve in Fig. 2(b), and its specific value is $0.9899$. It is observed that even for deviations of $|\delta T/T|\!=\!|\delta\Omega_{1}/\Omega_{1}|\!=\!|\delta\Omega_{2}/\Omega_{2}|\!=\!10\%$, the conversion efficiencies remain remarkably close to the ideal value. The results demonstrate that our protocol exhibits high robustness with respect to imperfections in operation.
@@
The above calculations demonstrate that the robustness, efficiency, and fast conversion of our protocol facilitate the preparation of ultracold tetratomic molecules. Nevertheless, it should be stressed that the resulting molecules are usually energetically unstable Kuznetsova et al. (2009), thereby making it highly necessary from a practical standpoint to rapidly transfer them to deeply-bound vibrational states Danzl et al. (2010); Kuznetsova et al. (2008); Mukherjee and Tomza (2026). To this end, we suggest a five-state C-STIREP technique, whose details will be further analyzed and presented in the following section.
@@ III.1 Model and Method
In this section, our goal is to generalize the STIREP into multi-state system, with the motivation to transfer the weakly-bound molecules into a sufficiently stable state. Inspired by the double-STIRAP transfer scheme Danzl et al. (2010); Kuznetsova et al. (2008), we restrict ourselves to a five-level chainwise-coupled molecular system. As presented in Fig. 6, all five molecular states are coupled in a distorted M-type configuration,
@@
in which four fields with Rabi frequencies $\Omega_{i}(i=1,2,3,4)$ are applied to move the molecules from a high vibrational level of the ground electronic state $|c\rangle$ (now relabeled as $|1\rangle$) to a sufficiently stable vibrational level of the ground electronic state $|5\rangle$. In the entire transfer process, three intermediate states are introduced to act as the “bridge”, where $|3\rangle$ is regarded as an intermediate vibrational level of ground electronic state, $|2\rangle$ and $|4\rangle$ represent two vibrational levels of the excited electronic state. To render this problem tractable, we shall analyze this problem under the collisionless limit, as done in prior Refs. Jing et al. (2008); Pu et al. (2007); Qian et al. (2010); Graefe et al. (2015). Accordingly, the Hamiltonian of the system can be described in the following form:
$\begin{aligned} \hat{\mathcal{H}}\!=\!\sum_{i=2}^{5}\Delta_{i}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{i}\!+\!\sum_{i=1}^{4}\left(\Omega_{i}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{i+1}\!+\!\mathrm{H.c.}\right).\end{aligned}$ (11)
@@
$\begin{aligned} i\dot{\psi}_{1}&\!=\!\Omega_{1}\psi_{2},\\ i\dot{\psi}_{2}&\!=\!\Omega_{1}\psi_{1}\!+\!\Delta\psi_{2}\!+\!\Omega_{2}\psi_{3},\\ i\dot{\psi}_{3}&\!=\!\Omega_{2}\psi_{2}\!+\!\Omega_{3}\psi_{4},\\ i\dot{\psi}_{4}&\!=\!\Omega_{3}\psi_{3}\!+\!\Delta\psi_{4}\!+\!\Omega_{4}\psi_{5},\\ i\dot{\psi}_{5}&\!=\!\Omega_{4}\psi_{4}.\end{aligned}$ (12)
@@
First, we postulate that $\Delta$ is the largest evolution frequency Liu et al. (2019b), $\Delta\!\gg\!\Omega_{i}(i\!=\!1,2,3,4)$. In this limit, adiabatic elimination (AE) enables the derivation of an effective three-state model,
@@
In which, the effective Rabi frequencies are $\omega_{13}=\!-\!\frac{\Omega_{1}\Omega_{2}}{\Delta}$ and $\omega_{35}=\!-\!\frac{\Omega_{3}\Omega_{4}}{\Delta}$; the effective detunings are $\varepsilon_{11}=\!-\!\frac{\Omega_{1}^{2}}{\Delta},\varepsilon_{22}=\!-\!\frac{\Omega_{2}^{2}+\Omega_{3}^{2}}{\Delta}$ and $\varepsilon_{33}=\!-\!\frac{\Omega_{4}^{2}}{\Delta}$, respectively. To continue, let us set $\varepsilon_{11}=\varepsilon_{22}=\varepsilon_{33}=\varepsilon_{0}$, from which the condition for generalized resonance locking follows,
@@
By further performing the simple transformation $\psi_{j}\!=\!\psi^{{}^{\prime}}_{j}e^{-i\varepsilon_{0}t}(j\!=\!1,3,5\!)$, we obtain
@@
$\begin{aligned} \omega_{13}\!&=-\!\frac{\Omega_{2}(t)\sqrt{\Omega_{2}^{2}(t)\!+\!\Omega_{3}^{2}(t)}}{\Delta},\\ \omega_{35}\!&=-\!\frac{\Omega_{3}(t)\sqrt{\Omega_{3}^{2}(t)\!+\!\Omega_{2}^{2}(t)}}{\Delta}.\end{aligned}$ (16)
@@
$\begin{aligned} &\psi^{{}^{\prime}}_{1}\!=\!\cos\phi(t)\cos\theta(t),\\ &\psi^{{}^{\prime}}_{3}\!=\!-i\sin\phi(t),\\ &\psi^{{}^{\prime}}_{5}\!=\!-\cos\phi(t)\sin\theta(t),\end{aligned}$ (17)
@@
$\begin{aligned} \omega_{13}&=\dot{\theta}(t)\cot\phi(t)\sin\theta(t)\!+\!\dot{\phi}(t)\cos\theta(t),\\ \omega_{35}&=\dot{\theta}(t)\cot\phi(t)\cos\theta(t)\!-\!\dot{\phi}(t)\sin\theta(t).\end{aligned}$ (18)
@@
Now we are ready to design the modified Rabi frequencies for the original five-state system. Like Eq. (16), we impose the following form on the two middle Rabi frequencies:
$\begin{aligned} \omega_{13}&=-\frac{\tilde{\Omega}_{2}(t)\sqrt{\tilde{\Omega}_{2}^{2}(t)\!+\!\tilde{\Omega}_{3}^{2}(t)}}{\Delta},\\ \omega_{35}&=-\frac{\tilde{\Omega}_{3}(t)\sqrt{\tilde{\Omega}_{2}^{2}(t)\!+\!\tilde{\Omega}_{3}^{2}(t)}}{\Delta}.\end{aligned}$ (19)
@@
$\begin{aligned} \tilde{\Omega}_{1}(t)=\tilde{\Omega}_{4}(t)=\sqrt{\tilde{\Omega}_{2}^{2}(t)+\tilde{\Omega}_{3}^{2}(t)}.\end{aligned}$ (20)
@@
$\begin{aligned} \tilde{\Omega}_{2}(t)&\!=\!\left(\frac{\Delta^{2}\omega^{4}_{13}}{\omega^{2}_{13}\!+\!\omega^{2}_{35}}\right)^{\frac{1}{4}},\\ \tilde{\Omega}_{3}(t)&\!=\!\left(\frac{\Delta^{2}\omega^{4}_{35}}{\omega^{2}_{13}\!+\!\omega^{2}_{35}}\right)^{\frac{1}{4}},\\ \tilde{\Omega}_{1}(t)&\!=\!\tilde{\Omega}_{4}(t)\!=\!\left[\Delta^{2}(\omega^{2}_{13}\!+\!\omega^{2}_{35})\right]^{\frac{1}{4}}.\end{aligned}$ (21)
With the newly designed Rabi frequencies, we thereby develop a five-state C-STIREP technique. It should be noted that, to ensure the validity of this method, the AE condition $\Delta\!\gg\!\tilde{\Omega}_{i}(t)(i\!=\!1,2,3,4)$ should still be satisfied Li and Chen (2016).
@@ III.2 Result and Discussion
Now we numerically investigate the coherent population dynamics of the C-STIREP method. Unlike Eq. (9) above, we demonstrate that the enhanced performance of C-STIREP can also be achieved by using other pulse shapes. As an example, we consider
@@
Once $\phi(t)$ and $\theta(t)$ are determined, the effective Rabi frequencies can be found according to Eq. (18).
@@
Figure. 7(a) illustrates the Rabi frequencies required to achieve the desired population transfer. From this figure, it can be observed that the first and fourth Rabi frequencies not only exhibit the same temporal profile and equal magnitudes, but also straddle the middle two. They exhibit profiles with sufficient smoothness, which thus poses no actual implementation challenges Du et al. (2016). The time evolution results from the initial state $|1\rangle$ to the target state $|5\rangle$ is displayed in Fig. 7(b). It can be observed that a complete state transfer can be achieved by employing four modified fields. A high transfer efficiency is helpful to minimize the loss in phase-space density when an ensemble of weakly-bound molecules is transferred to the ground state. In addition, we note that, throughout the entire transfer process, the intermediate state $|3\rangle$ receives some amount of transient population, while the populations of the two excited states are negligible. This process does not rely on the existence of dark states; instead, we utilize the AE protocol to achieve the stable decoupling of two excited states from dynamics. Thus, although these two states are prone to decaying out of the system, the transfer process remains scarcely swayed by them.
@@
Figure 8: (Color online) (a) The transient population of state $|3\rangle$ as a function of $\beta$, where the other parameters are consistent with those in Fig. 7; (b) The populations of state $|3\rangle$ and state $|5\rangle$ as a function of $t$ for different values of $T$, where $\beta$ is fixed at $0.1\pi$, and all other parameters are consistent with those in Fig. 7.
@@
Once a desired population transfer from $|1\rangle$ to $|5\rangle$ is achieved, we would like to impose further conditions of minimizing the amount of transient population that is in $|3\rangle$ throughout the process. Since the dynamics of the five-state system is equivalent to that of the reduced three-state one, the transient population of the intermediate state $|3\rangle$ can be accurately derived as $\sin^{2}\phi$. On this basis, we may attempt to reduce the value of $\phi$ to diminish the population of this state. We plot in Fig. 8(a) the transient population of state $|3\rangle$ as a function of $\beta$; please note that all other parameters are the same as those in Fig. 7. Clearly, the population of the intermediate state $|3\rangle$ witnesses a substantial decrease as the value of $\beta$ is gradually reduced. However, the population of this state never equals 0 since the limit of $\beta\rightarrow 0$ cannot be achieved in practical scenarios [see Eqs. (21) and (22)]. From the preceding analysis, it can be concluded that the five-state transfer scheme offers an additional benefit over the three-state counterpart: the former only requires us to control the transient population of the intermediate state $|3\rangle$, while the latter necessitates the control of the short-lived excited state. Given that the intermediate state $|3\rangle$ typically exhibits a sufficiently long lifetime, even a small amount of transient population within this state exerts a negligible influence on the final transfer efficiency; in contrast, the transfer efficiency of the three-state scheme is largely constrained by the characteristics of its intermediate short-lived excited state Ciamei et al. (2017); Li et al. (2017).
@@
Figure 9: (Color online) (a) Contour plot showing transfer efficiency versus $\delta\tilde{\Omega}_{1}$ and $\delta\tilde{\Omega}_{4}$, where $\tilde{\Omega}_{2}$ and $\tilde{\Omega}_{3}$ are set to their ideal values; (b) Contour plot showing transfer efficiency versus $\delta\tilde{\Omega}_{2}$ and $\delta\tilde{\Omega}_{3}$, where $\tilde{\Omega}_{1}$ and $\tilde{\Omega}_{4}$ passively vary according to condition (20). The parameter deviations are implemented on the basis of the adopted parameters in Fig. 8(b).
@@
Furthermore, we point out that the transient population of the intermediate state $|3\rangle$ can be further reduced by leveraging the speed advantage of C-STIREP. To verify this finding, we plot in Fig. 8(b) the variation of populations at time t=Tt\!=\!T of states $|3\rangle$ and $|5\rangle$ with the evolution time $T$. As can be seen from this figure, the proposed method is capable of significantly shortening the population transfer time without deteriorating the transfer efficiency. This interesting feature mitigates the interaction duration between the system and its surrounding environment, thereby enhancing the transfer efficiency for practical applications.
Finally, we have performed numerical simulations to verify the necessity of Condition (20). Figure. 9(a) shows a contour plot of transfer efficiency against the Rabi frequency deviations $\delta\tilde{\Omega}_{1}$ and $\delta\tilde{\Omega}_{4}$ when $\tilde{\Omega}_{2}$ and $\tilde{\Omega}_{3}$ are fixed at their ideal values. The transfer efficiency is defined as
$\begin{aligned} \digamma\!=\!\frac{P_{5,t\rightarrow\infty}}{P_{1,t\rightarrow 0}},\end{aligned}$ (23)
[…]